if a firm producing 100 units at $5.00 each experience an 80% experience curve, what will units costs be at 200 units of production?
Answer #1

Total cost for 100 units = 5*100 = $500

Now learning curve coefficient at N = 2 and 80% learning curve = 1.80 (total). This is as per learning curve coefficient table.

So total cost for 100*2 = 200 units (n = 2) will be = $500*1.80

= $900

Thus unit costs for the next 100 units (i.e. from 100 to 200) = (900-500/100) = $4 per unit

Unit costs for all the 200 units = $900/200 units = $4.5 per unit

Unit cost for next 100 units $ 4 per unit
Unit costs for all 200 units $4.5 per unit